given the linear equation y=3/7x-1, find the y coordinates of the points (-7),(0), (7) plot those points and graph the linear ewuation.
Answer by mananth(16949)

y=3/7x-1, find the y coordinates of the points (-7),(0), (7) plot those points
x=-7 y= 3/7 *-7 -1 = -4 (-7,-4)
x=0,y=-1 (0,-1)
x=7,y=2) (7,2)
Plot the points and draw the line graph